# Defining Cost Function of Steganography with Large Language Models

### v1

@@ Abstract
In this paper, we make the first attempt towards defining cost function of steganography with large language models (LLMs), which is totally different from previous works that rely heavily on expert knowledge or require large-scale datasets for cost learning. To achieve this goal, a two-stage strategy combining LLM-guided program synthesis with evolutionary search is applied in the proposed method. In the first stage, a certain number of cost functions in the form of computer program are synthesized from LLM responses to structured prompts. These cost functions are then evaluated with pretrained steganalysis models so that candidate cost functions suited to steganography can be collected. In the second stage, by retraining a steganalysis model for each candidate cost function, the optimal cost function(s) can be determined according to the detection accuracy. This two-stage strategy is performed by an iterative fashion so that the best cost function can be collected at the last iteration. Experiments show that the proposed method enables LLMs to design new cost functions of steganography that significantly outperform existing works in terms of resisting steganalysis tools, which verifies the superiority of the proposed method. To the best knowledge of the authors, this is the first work applying LLMs to the design of advanced cost function of steganography, which presents a novel perspective for steganography design and may shed light on further research.
@@ Introduction
There are two widely applied strategies for cost function definition in above minimum-distortion embedding framework. One is manually crafted, and the other is based on deep learning. The former relies heavily on the designer’s professional knowledge, whereas the latter learns the cost function through training a deep neural network, and therefore requires large-scale datasets. Generally, manually crafted methods [2, 3, 4] tend to utilize statistical features such as residual and context characteristics to construct a specific distortion function of image steganography. The advancement comes from the authors’ expertise in steganography. Deep learning based techniques, in contrast, rely relatively less on human experience, but requires a large amount of data for training a neural network that can generate a cost map or a probability map used to define the steganographic cost. Many advanced methods have been proposed in this direction. In summary, the evolution of cost function design has progressed through two distinct generations: experience-based and learning-based methods.
@@
The objective of the Stage 1 is to efficiently identify and preliminarily filter a set of candidate cost functions, ensuring both the scalability and search efficiency of the framework. To achieve this objective, an iterative evolution-based search strategy driven by LLM is employed. With structured prompts, the LLM iteratively refines existing cost function to produce multiple improved variants. These variants then undergo rapid evaluation, and those that execute successfully and yield valid scores are incorporated into the program database as candidate cost functions for steganography. All cost functions are in the form of computer program. We describe the workflow of Stage 1 in the following.
@@
In Stage 2, the threshold for the number of programs selected for accurate evaluation is set as 10, meanng that if the total number of programs in PS is less than 10, all the programs will undergo accurate evaluation, otherwise 50% programs are randomly selected for accurate evaluation. An optimized program is designated only if its accurate evaluation score is no less than the score of the initial program, i.e., $s_{\text{init}}$, by a margin $\epsilon=1\%$. For the potential candidate function, to ensure the reliability of the performance improvement, we convert the evolved Python programs into MATLAB implementations. These MATLAB programs are then subjected to a rigorous multi-rate assessment at 0.4 bpp, 0.3 bpp, 0.2 bpp, and 0.1 bpp, and compared against their corresponding initial programs. We hope to generate three new cost functions outperforming WOW, HILL and SUNIWARD, respectively.
@@
The cost functions evolved from WOW, HILL, and SUNIWARD are referred to as the WOW-evolved, HILL-evolved, and SUNIWARD-evolved cost functions, respectively. Figure 2 illustrates the embedding change probabilities of a cover image under a relative payload of 0.4 bpp using WOW, WOW-evolved, HILL, HILL-evolved, S-UNIWARD-evolved and S-UNIWARD-evolved cost functions. The probability map of modification is encoded in brightness, where brighter areas represent lower probabilities of embedding. The presented cover image contains numerous horizontal and vertical edges, along with textured regions. For clarity, a cropped and enlarged portion of the image is displayed.
@@
We report the performance of the evolved algorithms against steganalysis. The performance is evaluated across various embedding rates ranging from 0.1 bpp to 0.4 bpp with a step of 0.1. The detection error rate in Eq. (13) is adopted as the indicator. The experimental setting is the same as the process of evolutionary, e.g., SRNet is used as the steganalyzer, and the ratio between training set, validation set and test set is 5:1:4. The experimental results shown in Table 1 indicates that all evolved algorithms consistently outperform their original counterparts, exhibiting enhanced security across every embedding strength and improved resistance to steganalysis. To quantify this improvement, we determine the (relative) performance gain, defined as $\Delta^{\uparrow}=(P_{\text{E}}^{\star}-P_{\text{E}}^{\circ})/P_{\text{E}}^{\circ}$, where $P_{\text{E}}^{\star}$ and $P_{\text{E}}^{\circ}$ correspond to the evolved algorithm and raw version.

### v2

@@ Abstract
In this paper, we make the first attempt towards defining cost function of steganography with large language models (LLMs), which is totally different from previous works that rely heavily on expert knowledge or require large-scale datasets for cost learning. To achieve this goal, a two-stage strategy combining LLM-guided program synthesis with evolutionary search is applied in the proposed method. In the first stage, a certain number of cost functions in the form of computer programs are synthesized from LLM responses to structured prompts. These cost functions are then evaluated with pretrained steganalysis models so that candidate cost functions suited to steganography can be collected. In the second stage, by retraining a steganalysis model for each candidate cost function, the optimal cost function(s) can be determined according to the detection accuracy. This two-stage strategy is performed by an iterative fashion so that the best cost function can be collected at the last iteration. Experiments show that the proposed method enables LLMs to design new cost functions of steganography that significantly outperform existing works in terms of resisting steganalysis tools, which verifies the superiority of the proposed method. To the best knowledge of the authors, this is the first work applying LLMs to the design of advanced cost function of steganography, which presents a novel perspective for steganography design and may shed light on further research.
@@ Introduction
There are two widely applied strategies for cost function definition in above minimum-distortion embedding framework. One is manually crafted, and the other is based on deep learning. The former relies heavily on the designer’s professional knowledge, whereas the latter learns the cost function through training a deep neural network, and therefore requires large-scale datasets. Generally, manually crafted methods [2, 3, 4] tend to utilize statistical features such as residual and context characteristics to construct a specific distortion function of image steganography. The advancement comes from the authors’ expertise in steganography. Deep learning based techniques, in contrast, rely relatively less on human experience, but require a large amount of data for training a neural network that can generate a cost map or a probability map used to define the steganographic cost. Many advanced methods have been proposed in this direction. In summary, the evolution of cost function design has progressed through two distinct generations: experience-based and learning-based methods.
@@
The objective of the Stage 1 is to efficiently identify and preliminarily filter a set of candidate cost functions, ensuring both the scalability and search efficiency of the framework. To achieve this objective, an iterative evolution-based search strategy driven by LLM is employed. With structured prompts, the LLM iteratively refines existing cost functions to produce multiple improved variants. These variants then undergo rapid evaluation, and those that execute successfully and yield valid scores are incorporated into the program database as candidate cost functions for steganography. All cost functions are in the form of computer programs. We describe the workflow of Stage 1 in the following.
@@
In Stage 2, the threshold for the number of programs selected for accurate evaluation is set as 10, meaning that if the total number of programs in PS is less than 10, all the programs will undergo accurate evaluation, otherwise 50% programs are randomly selected for accurate evaluation. An optimized program is designated only if its accurate evaluation score is no less than the score of the initial program, i.e., $s_{\text{init}}$, by a margin $\epsilon=1\%$. For the potential candidate function, to ensure the reliability of the performance improvement, we convert the evolved Python programs into MATLAB implementations. These MATLAB programs are then subjected to a rigorous multi-rate assessment at 0.4 bpp, 0.3 bpp, 0.2 bpp, and 0.1 bpp, and compared against their corresponding initial programs. We hope to generate three new cost functions outperforming WOW, HILL and SUNIWARD, respectively.
@@
The cost functions evolved from WOW, HILL, and SUNIWARD are referred to as the WOW-evolved, HILL-evolved, and SUNIWARD-evolved cost functions, respectively. Figure 2 illustrates the embedding change probabilities of a cover image under a relative payload of 0.4 bpp using WOW, WOW-evolved, HILL, HILL-evolved, SUNIWARD and SUNIWARD-evolved cost functions. The probability map of modification is encoded in brightness, where brighter areas represent lower probabilities of embedding. The presented cover image contains numerous horizontal and vertical edges, along with textured regions. For clarity, a cropped and enlarged portion of the image is displayed.
@@
We report the performance of the evolved algorithms against steganalysis. The performance is evaluated across various embedding rates ranging from 0.1 bpp to 0.4 bpp with a step of 0.1. The detection error rate in Eq. (13) is adopted as the indicator. The experimental setting is the same as the process of evolutionary, e.g., SRNet is used as the steganalyzer, and the ratio between training set, validation set and test set is 5:1:4. The experimental results shown in Table 1 indicate that all evolved algorithms consistently outperform their original counterparts, exhibiting enhanced security across every embedding strength and improved resistance to steganalysis. To quantify this improvement, we determine the (relative) performance gain, defined as $\Delta^{\uparrow}=(P_{\text{E}}^{\star}-P_{\text{E}}^{\circ})/P_{\text{E}}^{\circ}$, where $P_{\text{E}}^{\star}$ and $P_{\text{E}}^{\circ}$ correspond to the evolved algorithm and raw version.